# Async Learned User Embeddings for Ads Delivery Optimization

### v1

@@ Abstract.
User representation is crucial for recommendation systems as it helps to deliver personalized recommendations by capturing user preferences and behaviors in low-dimensional vectors. High-quality user embeddings can capture subtle preferences, enable precise similarity calculations, and adapt to changing preferences over time to maintain relevance. The effectiveness of recommendation systems depends significantly on the quality of user embedding. We propose to asynchronously learn high fidelity user embeddings for billions of users each day from sequence based multimodal user activities in Meta platforms through a Transformer-like large scale feature learning module. The async learned user representations embeddings (ALURE) are further converted to user similarity graphs through graph learning and then combined with user realtime activities to retrieval highly related ads candidates for the entire ads delivery system. Our method shows significant gains in both offline and online experiments.
@@ 1. Introduction
In this paper, we introduce how to generate high fidelity compact user representations and leverage it to improve retrieval performance. In Meta, we use a multi-stage ads delivery system (Wang et al., 2023; Covington et al., 2016) as shown in Figure 1.
@@
Figure 2. The feedback loop for User Representation Learning in Meta’s ads delivery system

### v2

@@ Abstract.
In recommendation systems, high-quality user embeddings can capture subtle preferences, enable precise similarity calculations, and adapt to changing preferences over time to maintain relevance. The effectiveness of recommendation systems depends on the quality of user embedding. We propose to asynchronously learn high fidelity user embeddings for billions of users each day from sequence based multimodal user activities through a Transformer-like large scale feature learning module. The async learned user representations embeddings (ALURE) are further converted to user similarity graphs through graph learning and then combined with user realtime activities to retrieval highly related ads candidates for the ads delivery system. Our method shows significant gains in both offline and online experiments.
@@ 1. Introduction
In this paper, we introduce how to generate high fidelity compact user representations and leverage it to improve retrieval performance. In industry, ads are delivered through a multi-stage ranking system (Wang et al., 2023; Covington et al., 2016) as shown in Figure 1.
@@
Figure 2. The feedback loop for User Representation Learning